• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Boeglin

    Log entry time 16:04:37 on June 6,2010

    Entry number 324893

    keyword=Shift Summary

    Do Harp scan: beam should be 150 < sigma < 250
    Start long recovery
    9:00 MCC called, they can deliver some beam but they have to replace a piece if interlock electronics which will take another access.

    9:20 got beam into the Hall: Harp Scan: 254 x 201 microns
    9:50 checked beam size and position with BeO and Pb #3
    observed a really low rate on Pb
    9:55 work continues at Accelecrator to replace electronics
    10:30 set laser att. to 55 with slit 17
    took spot_L & spot_R on BeO
    move to Pb #3: do spot_L & spot_R, looked OK -> ramp up to 50uA
    10:55 after many beam trips MCC informed that they get injector BLM trips above 45uA -> running at 40uA
    encountered some problems with DAQ
    first production runs : 4335. Some cuts need to be adjusted in order
    for the online analysis to work. We can run at the moment with 40-42uA only

    4336 should be a good run at 40uA
    try 46uA
    4337 HV trip: compton and left arm detectors ended run
    4338 restored HV run at 42uA
    14:05 MCC called top get 1/2 hour of beam to work on the BLM trip problem in order ot be anble to get 50uA -> ended run
    2:30 got beam back at 50uA start new run (while ramping)
    2:27 started run 4339
    3:20 start run 4340 first run with constant 50uA
    3:45 Right Q1 trip, no obvious reason, reset it. did NOT stop run 4340